OpenStudy (anonymous):

what would be the limit of ((1-x^3)/x^2+9x))^5 as x goes to negative infinity?

OpenStudy (aum):

Is this the correct problem? \[ \lim_{x \rightarrow -\infty}\left( \frac{1-x^3}{x^2+9x} \right )^5 \]

OpenStudy (anonymous):

yes

OpenStudy (aum):

\[ \lim_{x \rightarrow -\infty}\left( \frac{1-x^3}{x^2+9x} \right )^5 = \lim_{x \rightarrow -\infty}\left( = \frac{x^2(1/x^2-x)}{x^2(1+9/x)} \right )^5 \\ \lim_{x \rightarrow -\infty}\left( \frac{((1/x^2-x)}{(1+9/x)} \right )^5 = \infty\\ \]

OpenStudy (anonymous):

ohh okay thank you (:

OpenStudy (aum):

you are welcome.